How do you do a problem like this:
On Marikas 8th birth day in 2004, her father said, "My age is now four times yours." In what year will her father be able to say my age is 3 times yours?

Answer by mananth(16946)

marika's age = 8 years
Father's age 4*8=32 years
let him be 3 times her age after x years
Marika's age after x years 8+x
Father's age after x years 32+x
32+x=3(8+x)
32+x=24+3x
3x-x=32-24
2x=8
/2
x=4
after 4 years